Hello. I would like to know how to write a proof for this problem:
Given: AC=AB, AC=4x+1, AB=6x-13
Prove: x=7

Answer by mananth(16949)

AC=AB, AC=4x+1, AB=6x-13
4x+1=6x-13
6x-4x=13+1
2x=14
/2
x=7